# Properties of leveled spatial graphs

### v1

@@ 5 Relation between leveled and paneled spatial graphs
A paneled embedding of a spatial graph $\mathcal{G}$ is not necessarily semi-leveled. Indeed, consider two copies of $K_{3,3}$ connected by an edge, as in fig. 6. The embedding is paneled, because $K_{3,3}$ is paneled, as shown by Robertson, Seymour and Thomas ([8]). However, the embedding is not semi-leveled, because at least one fragment with respect to any cycle is non-planar.
@@
If the fragment $f_{i}$ in the sequence $f_{1},\dots,f_{t}$ crosses under the fragment $f_{k}$, with $k>i+1$, then we can reduce the sequence by removing all fragments from $f_{i+1}$ to $f_{k-1}$. In this way, we obtain a sequence $f_{1}<f_{2}<\dots<f_{t}<f_{1}$ where $f_{i}$ does not cross under any fragment other than $f_{i+1}$. In the same way, $f_{t}$ crosses under only $f_{1}$. This also implies that $f_{i}$ crosses over only $f_{i-1}$ for $i=2,\dots,t$ and $f_{1}$ crosses over only $f_{t}$. Given that fragments are connected, we can restrict the sequence of fragments to sequence of edges $e_{1},\dots e_{t}$ with $e_{1}<\dots<e_{t}<e_{1}$, where $e_{i}$ crosses under only $e_{i+1}$ (modulo $t$).
We now prove that the subdiagram $\tilde{D}\subset D$ that is formed by the spine of $D$ and the edges of the sequence $e_{1},\dots e_{t}$ from above depicts a nontrivial knot or link for $t\geq 4$, implying that $\mathcal{G}$ cannot be paneled. We deal with the case $t=3$ later.
@@
Suppose that $e_{i-1}$ crosses under $e_{i}$ at least twice. This means that there exists at least one region in the diagram $\tilde{D}$ formed by a superarc of $e_{i}$ and a subarc of $e_{i-1}$ (two such regions are highlighted with a red circle in fig. 7). In order for the diagram to be reduced, another edge $e$ would need to cross this region such that $e$ crosses over $e_{i}$ and under $e_{i-1}$, since otherwise a Reidemeister move of type 2 can be performed to reduce the crossings between $e_{i}$ and $e_{i-1}$. Since $e_{i-1}$ only has crossings with $e_{i}$ and $e_{i-2}$, $e$ must equal $e_{i-2}$. Also, since $e_{i}$ only has crossings with $e_{i-1}$ and $e_{i+1}$, $e$ must equal $e_{i+1}$. Since $t\geq 4$, $e_{i-2}\neq e_{i+1}$ and no edge $e$ exists. Therefore, a Reidemeister move of type 2 can be performed to remove the region. Applying the same reasoning to every region formed between $e_{i}$ and $e_{i-1}$ concludes that these fragments cross each other exactly once.
Thus, if $t\geq 4$, $e_{i}$ crosses under $e_{i+1}$ exactly once. Up to mirror image, this means that we obtain an embedding as in fig. 8. In such an embedding, there exists a torus link or a torus knot $T(t,2)$. Indeed, consider the cycle obtained by starting at edge $e_{1}$, running through edge $e_{3}$, $e_{5}$ and so on, as depicted in blue in fig. 8 for $6$ edges. If the number $t$ of edges is odd, this process ends with the torus knot $T(t,2)$. If $t$ is even, we obtain the torus link $T(t,2)$ by starting another cycle at edge $e_{2}$ and running through even-indexed edges. In both cases, the spatial graph contains a link or a knot and hence it cannot be paneled.
@@
We are left with the case $t=3$. For $e_{1}<e_{2}<e_{3}<e_{1}$, the graph $\tilde{\mathcal{G}}\subset\mathcal{G}$ that is determined by $\tilde{D}$ has 6 vertices and 9 edges. According to Kuratowski theorem, such a graph is not planar if and only if it is isomorphic to $K_{3,3}$ ([6]). Suppose that the underlying graph $\tilde{G}$ of $\tilde{\mathcal{G}}$ is not $K_{3,3}$ and therefore is planar. As $\tilde{\mathcal{G}}$ is not leveled due to the existence of the edge sequence, this implies that it cannot be trivial, as every trivial planar graph is leveled by considering the cycle in the boundary of any face as spine. Since by Theorem 2.2 of [8] an embedding of a planar graph is paneled if and only if it is trivial, it follows that $\mathcal{G}$ is not paneled. It remains to show that the underlying graph $\tilde{G}$ of $\tilde{\mathcal{G}}$ cannot be $K_{3,3}$. We do this by arguing that every semi-leveled and paneled embedding of $K_{3,3}$ is leveled: Theorem (3.1) of [8] states that there exist only two non-ambient-isotopic paneled embeddings of $K_{3,3}$, depicted in fig. 9. Since both these embeddings of $K_{3,3}$ are leveled, it follows that every paneled and semi-leveled embedding of $K_{3,3}$ is leveled.
@@
To prove this result, we use the following lemma of Böhme ([4]).
@@ 6 A new invariant to measure non-planarity
As a spatial graph is defined up to ambient isotopy, also the level number of a spatial graph should be defined up to ambient isotopy. Since a leveled embedding might have more than one cycle that can act as spine, and considering another spine of a spatial graph can result in a different number of levels (see for example fig. 10, where two different choices of the spine for a hamiltonian leveled spatial graph have different number of levels), we need to range over all possible spines for a leveled embedding of the spatial graph.
@@
Consider the hamiltonian graph $K_{4,4}$. Both its level number $\ell(K_{4,4})$ and its hamiltonian level number $h\ell(K_{4,4})$ equal 4, as we show in theorem 3. Consider instead the hamiltonian graph $G$ obtained from $K_{4,4}$ by adding the four dashed edges in fig. 12. One might expect $h\ell(G)\geq h\ell(K_{4,4})$ as $K_{4,4}$ is a subgraph of $G$ obtained by removing some edges. However, $h\ell(G)=3$ as shown in fig. 12, and since $\ell(G)\leq h\ell(G)$, it also holds that $\ell(G)\leq\ell(K_{4,4})$.
@@ 6.1 Comparison to existing similar invariants
Lastly, we show that $h\ell(K_{n})=\ell(K_{n}).$ Since $\ell(G)\leq h\ell(G)$, it is sufficient to show that considering a non-hamiltonian spine for $K_{n}$ does not yield a level number smaller than $\lceil\frac{n}{2}\rceil$. If the spine was a non-hamiltonian cycle and there were two or more vertices of $K_{n}$ not contained in this cycle, the vertices that are not part of the cycle must belong to the same level given that they are adjacent. In that case, both vertices $v$ and $w$ and all edges issuing from them need to be embedded in a disk $D$ whose boundary coincides with the spine. In $D$ we can first place $v$ and connect it with all vertices on the spine, decomposing $D$ into triangles. Then, we would need to place $w$ into a triangle, meaning that we can connect it with at most three vertices. If $n\geq 5$, $w$ has degree at least 4 and therefore $w$ and $v$ cannot belong to the same level. Therefore, we are left to consider spines with at least $n-1$ vertices. If exactly one vertex does not belong to the spine, we obtain the leveled embedding described in the proof of proposition 5. As for this embedding the inequality of proposition 5 is sharp, its level number equals $h\ell(K_{n-1})+1$. As $h\ell(K_{n-1})+1=\lceil\frac{n-1}{2}\rceil+1\geq\lceil\frac{n}{2}\rceil$ for $n\geq 5$, the statement follows.∎
@@
$h\ell(K_{n})=\begin{cases}0,&\text{if}\ n=3\\ \lceil\frac{n}{2}\rceil,&\text{if}\ n\geq 4.\end{cases}.$
@@
    We now prove that $h\ell(K_{n,n})=n.$ Consider the vertices of $K_{n,n}$ to be on a cycle and label them $x_{1},x_{2},\dots,x_{2n}$, such that $x_{i}\in A$ if $i$ is odd and $x_{i}\in B$ if $i$ is even. All edges issuing from vertex $x_{2i+1}\in A$ can be assigned to the same level, and this holds for $i=0,\dots,n-1$, implying that $h\ell(K_{n,n})\leq n.$ We are left to show that $n$ levels are needed. If $n$ is odd, the edges $(x_{1},x_{n+1}),(x_{3},x_{n+3}),\dots,(x_{2n-1},x_{n-1})$ all conflict with each other. Therefore, there must be at least $n$ levels. If $n$ is even, the edges $(x_{1},x_{n}),(x_{2},x_{n+1}),\dots,(x_{n-1},x_{2n-2})$ all conflict with each other. Therefore, there must be at least $n-1$ levels. Assign edge $(x_{i},x_{i+n-1})$ to level $l_{i}$ for $i=1,\dots,n-1$. The edge $e=(x_{n},x_{2n-1})$ conflicts with the $n-1$ previous edges $(x_{2},x_{n+1}),\dots,(x_{n-1},x_{2n-2})$ but not with $(x_{1},x_{n})$. If we assign $e$ to level $l_{n}$, the proof is complete because $K_{n,n}$ has at least $n$ levels. Otherwise, we have to assign $e$ to level $l_{1}$. The edge $(x_{n+1},x_{2n})$ is assigned to level $l_{2}$ since it conflicts with $(x_{3},x_{n+2}),\dots,(x_{n},x_{2n-1})$. We continue in this fashion, assigning edge $(x_{n+t},x_{2n+t-1})$ to level $l_{t+1}$ for $t=0,\dots,n-2$, where the indices need to be considered modulo $2n$. Consider now the edge $(x_{2n-1},x_{3n-2})=(x_{2n-1},x_{n-2}).$ It conflicts with the $n-1$ edges $(x_{n+1},x_{2n}),\dots,(x_{2n-2},x_{3n-3})=(x_{2n-2},x_{n-3})$, which are assigned to levels $l_{2},\dots,l_{n-1}$. It also conflicts with edge $(x_{1},x_{n})$, assigned to level $l_{1}$. Therefore, at least $n$ levels are necessary.
  • If $m>n$, $K_{m,n}$ has no hamiltonian cycle. If $n>2$, an argument similar to the one showing that $\ell(K_{n,n})=h\ell(K_{n,n})$ ensures that the $m-n$ vertices that do not belong to the spine must be assigned to different levels, and that no edge other than those issuing from each of those vertices can be assigned to the same level. Therefore, $\ell(K_{m,n})=(m-n)+h\ell(K_{n,n})=m$.

### v2

@@ 5 Relation between leveled and paneled spatial graphs
A paneled embedding of a spatial graph $\mathcal{G}$ is not necessarily semi-leveled. Indeed, consider two copies of $K_{3,3}$ connected by an edge, as in Figure 7. The embedding is paneled, because $K_{3,3}$ is paneled, as shown by Robertson, Seymour and Thomas ([10]). However, the embedding is not semi-leveled, because at least one fragment with respect to any cycle is non-planar.
@@
If the fragment $f_{i}$ in the sequence $f_{1},\dots,f_{t}$ crosses under the fragment $f_{k}$, with $k>i+1$, then we can reduce the sequence by removing all fragments from $f_{i+1}$ to $f_{k-1}$. In this way, we obtain a sequence $f_{1}<f_{2}<\dots<f_{t}<f_{1}$ where $f_{i}$ does not cross under any fragment other than $f_{i+1}$. In the same way, $f_{t}$ crosses under only $f_{1}$. This also implies that $f_{i}$ crosses over only $f_{i-1}$ for $i=2,\dots,t$ and $f_{1}$ crosses over only $f_{t}$. Given that fragments are connected, we can restrict the sequence of fragments to sequence of edges $e_{1},\dots,e_{t}$ with $e_{1}<\dots<e_{t}<e_{1}$, where $e_{i}$ crosses under only $e_{i+1}$ (modulo $t$) and where $e_{a}<e_{b}$ if the edge $e_{a}$ crosses under edge $e_{b}$ non-trivially (meaning that no Reidemeister moves of type 2 can be applied to remove the crossing between $e_{a}$ and $e_{b}$).
We now prove that the subdiagram $\tilde{D}\subset D$ that is formed by the spine of $D$ and the edges of the sequence $e_{1},\dots,e_{t}$ from above depicts a nontrivial knot or link for $t\geq 4$, implying that $\mathcal{G}$ cannot be paneled. We deal with the case $t=3$ later.
@@
Suppose that $e_{i-1}$ crosses under $e_{i}$ at least twice. This means that there exists at least one region in the diagram $\tilde{D}$ formed by a superarc of $e_{i}$ and a subarc of $e_{i-1}$ (two such regions are highlighted with a red circle in Figure 8). In order for the diagram to be reduced, another edge $e$ would need to cross this region such that $e$ crosses over $e_{i}$ and under $e_{i-1}$, since otherwise a Reidemeister move of type 2 can be performed to reduce the crossings between $e_{i}$ and $e_{i-1}$. Since $e_{i-1}$ only has crossings with $e_{i}$ and $_{i-2}, ee$ must equal $e_{i-2}$. Also, since $e_{i}$ only has crossings with $e_{i-1}$ and $_{i+1}, ee$ must equal $e_{i+1}$. Since $t\geq 4$, $e_{i-2}\neq e_{i+1}$ and no edge $e$ exists. Therefore, a Reidemeister move of type 2 can be performed to remove the region. Applying the same reasoning to every region formed between $e_{i}$ and $e_{i-1}$ concludes that these fragments cross each other exactly once.
Thus, if $t\geq 4$, $e_{i}$ crosses under $e_{i+1}$ exactly once. Up to mirror image, this means that we obtain an embedding as in Figure 9. In such an embedding, there exists a torus link or a torus knot $T(t,2)$. Indeed, consider the cycle obtained by starting at edge $e_{1}$, running through edge $e_{3}$, $e_{5}$ and so on, as depicted in blue in Figure 9 for $6$ edges. If the number $t$ of edges is odd, this process ends with the torus knot $T(t,2)$. If $t$ is even, we obtain the torus link $T(t,2)$ by starting another cycle at edge $e_{2}$ and running through even-indexed edges. In both cases, the spatial graph contains a link or a knot and hence it cannot be paneled.
@@
We are left with the case $t=3$. For $e_{1}<e_{2}<e_{3}<e_{1}$, the graph $\tilde{\mathcal{G}}\subset\mathcal{G}$ that is determined by $\tilde{D}$ has 6 vertices and 9 edges. According to Kuratowski theorem, such a graph is not planar if and only if it is isomorphic to $K_{3,3}$ ([7]). Suppose that the underlying graph $\tilde{G}$ of $\tilde{\mathcal{G}}$ is not $K_{3,3}$ and therefore is planar. As $\tilde{\mathcal{G}}$ is not leveled due to the existence of the edge sequence, this implies that it cannot be trivial, as every trivial planar graph is leveled by considering the cycle in the boundary of any face as spine. Since by Theorem 2.2 of [10] an embedding of a planar graph is paneled if and only if it is trivial, it follows that $\mathcal{G}$ is not paneled. It remains to show that the underlying graph $\tilde{G}$ of $\tilde{\mathcal{G}}$ cannot be $K_{3,3}$. We do this by arguing that every semi-leveled and paneled embedding of $K_{3,3}$ is leveled: Theorem (3.1) of [10] states that there exist only two non-ambient-isotopic paneled embeddings of $K_{3,3}$, depicted in Figure 10. Since both these embeddings of $K_{3,3}$ are leveled, it follows that every paneled and semi-leveled embedding of $K_{3,3}$ is leveled.
@@
To prove this result, we use the following lemma of Böhme ([4]). A circuit is a closed path in a graph where edges cannot be repeated.
@@ 6 A new invariant to measure non-planarity
As a spatial graph is defined up to ambient isotopy, also the level number of a spatial graph should be defined up to ambient isotopy. Since a leveled embedding might have more than one cycle that can act as spine, and considering another spine of a spatial graph can result in a different number of levels (see for example Figure 11, where two different choices of the spine for a Hamiltonian leveled spatial graph have different number of levels), we need to range over all possible spines for a leveled embedding of the spatial graph.
@@
Consider the Hamiltonian graph $K_{4,4}$. Both its level number $\ell(K_{4,4})$ and its Hamiltonian level number $h\ell(K_{4,4})$ equal 4, as we show in theorem 3. Consider instead the Hamiltonian graph $G$ obtained from $K_{4,4}$ by adding the four dashed edges in Figure 13. One might expect $h\ell(G)\geq h\ell(K_{4,4})$ as $K_{4,4}$ is a subgraph of $G$ obtained by removing some edges. However, $h\ell(G)=3$ as shown in Figure 13, and since $\ell(G)\leq h\ell(G)$, it also holds that $\ell(G)\leq\ell(K_{4,4})$.
@@ 6.1 Comparison to existing similar invariants
Lastly, we show that $h\ell(K_{n})=\ell(K_{n}).$ Since $\ell(G)\leq h\ell(G)$, it is sufficient to show that considering a non-Hamiltonian spine for $K_{n}$ does not yield a level number smaller than $\lceil\frac{n}{2}\rceil$. If the spine was a non-Hamiltonian cycle and there were two or more vertices of $K_{n}$ not contained in this cycle, the vertices that are not part of the cycle must belong to the same level given that they are adjacent. In that case, both vertices $v$ and $w$ and all edges issuing from them need to be embedded in a disk $D$ whose boundary coincides with the spine. In $D$ we can first place $v$ and connect it with all vertices on the spine, decomposing $D$ into triangles. Then, we would need to place $w$ into a triangle, meaning that we can connect it with at most three vertices. If $n\geq 5$, $w$ has degree at least 4 and therefore $w$ and $v$ cannot belong to the same level. Therefore, we are left to consider spines with at least n1n-1 vertices. If exactly one vertex does not belong to the spine, we obtain the leveled embedding described in the proof of proposition 5. As for this embedding the inequality of proposition 5 is sharp, its level number equals $h\ell(K_{n-1})+1$. As $h\ell(K_{n-1})+1=\lceil\frac{n-1}{2}\rceil+1\geq\lceil\frac{n}{2}\rceil$ for $n\geq 5$, the statement follows.∎
@@
$h\ell(K_{n})=\begin{cases}0,&\text{if}\ n=3\\ \lceil\frac{n}{2}\rceil,&\text{if}\ n\geq 4.\end{cases}$
@@
    We now prove that $h\ell(K_{n,n})=n.$ Consider the vertices of $K_{n,n}$ to be on a cycle and label them $x_{1},x_{2},\dots,x_{2n}$, such that $x_{i}\in A$ if $i$ is odd and $x_{i}\in B$ if $i$ is even. All edges issuing from vertex $x_{2i+1}\in A$ can be assigned to the same level, and this holds for $i=0,\dots,n-1$, implying that $h\ell(K_{n,n})\leq n.$ We are left to show that $n$ levels are needed. If $n$ is odd, the edges $(x_{1},x_{n+1}),(x_{3},x_{n+3}),\dots,(x_{2n-1},x_{n-1})$ all conflict with each other. Therefore, there must be at least $n$ levels. If $n$ is even, the edges $(x_{1},x_{n}),(x_{2},x_{n+1}),\dots,(x_{n-1},x_{2n-2})$ all conflict with each other. Therefore, there must be at least n1n-1 levels. Assign edge $(x_{i},x_{i+n-1})$ to level $l_{i}$ for $i=1,\dots,n-1$. The edge $e=(x_{n},x_{2n-1})$ conflicts with the n1n-1 previous edges $(x_{2},x_{n+1}),\dots,(x_{n-1},x_{2n-2})$ but not with $(x_{1},x_{n})$. If we assign $e$ to level $l_{n}$, the proof is complete because $K_{n,n}$ has at least $n$ levels. Otherwise, we have to assign $e$ to level $l_{1}$. The edge $(x_{n+1},x_{2n})$ is assigned to level $l_{2}$ since it conflicts with $(x_{3},x_{n+2}),\dots,(x_{n},x_{2n-1})$. We continue in this fashion, assigning edge $(x_{n+t},x_{2n+t-1})$ to level $l_{t+1}$ for $t=0,\dots,n-2$, where the indices need to be considered modulo $2n$. Consider now the edge $(x_{2n-1},x_{3n-2})=(x_{2n-1},x_{n-2}).$ It conflicts with the n1n-1 edges $(x_{n+1},x_{2n}),\dots,(x_{2n-2},x_{3n-3})=(x_{2n-2},x_{n-3})$, which are assigned to levels $l_{2},\dots,l_{n-1}$. It also conflicts with edge $(x_{1},x_{n})$, assigned to level $l_{1}$. Therefore, at least $n$ levels are necessary.
  • If $m>n$, $K_{m,n}$ has no Hamiltonian cycle. If $n>2$, an argument similar to the one showing that $\ell(K_{n,n})=h\ell(K_{n,n})$ ensures that the mnm-n vertices that do not belong to the spine must be assigned to different levels, and that no edge other than those issuing from each of those vertices can be assigned to the same level. Therefore, $\ell(K_{m,n})=(m-n)+h\ell(K_{n,n})=m$.